OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 2r2 + 3s3 - r2 + 4t2 - r2 if r = -2, s = -3, and t = 5. -19 -37 27 19

OpenStudy (anonymous):

Also can you help with this one? Evaluate the following expression using the values given: Find 3x2 - y3 - y3 - z if x = 3, y = -2, and z = -5. Numerical Answers Expected!

OpenStudy (anonymous):

\[2(-2)^2+3(-3)^3-(-2)^2+4(5)^2-(-2)^2=19\]

OpenStudy (anonymous):

\[3(3)^2-(-2)^3-(-2)^3-(-5)=48\]